### v1

# Yuxin Feng4, Long Ma2†, Bo Du5, Pheng-Ann Heng1

Abstract
In laparoscopic surgery, a clear and high-quality visual field is critical for surgeons to make accurate intraoperative decisions. However, persistent visual degradation, including smoke generated by energy devices, lens fogging from thermal gradients, and lens contamination due to blood or tissue fluid splashes during surgical procedures, severely impair visual clarity. These degenerations can seriously hinder surgical workflow and pose risks to patient safety. To systematically investigate and address various forms of surgical scene degradation, we introduce a real-world open-source surgical image restoration dataset covering laparoscopic environments, called SurgClean, which involves multi-type image restoration tasks, e.g., desmoking, defogging, and desplashing. SurgClean comprises 1,020 images with diverse degradation types and corresponding paired reference labels. Based on SurgClean, we establish a standardized evaluation benchmark and provide performance for 22 representative generic task-specific image restoration approaches, including 12 generic and 10 task-specific image restoration approaches. Experimental results reveal substantial performance gaps relative to clinical requirements, highlighting a critical opportunity for algorithm advancements in intelligent surgical restoration. Furthermore, we explore the degradation discrepancies between surgical and natural scenes from structural perception and semantic understanding perspectives, providing fundamental insights for domain-specific image restoration research. Our work aims to empower the capabilities of restoration algorithms to increase surgical environments and improve the efficiency of clinical procedures.
1 Introduction
Backgrounds: In clinical minimally invasive surgery, a clear surgical field of view and improved image quality in the operative area are critical for preventing unexpected complications and ensuring surgical safety. Especially in narrow laparoscopic environments, visual clarity could be severely impaired by various intraoperative factors, e.g., smoke generated by energy instruments, fog formation on lenses due to temperature differences, and lens contamination caused by splashes of blood or tissue fluids during surgical manipulations, as illustrated in Fig. 1. Diverse types of degradation seriously hinder surgeons’ observation and judgment of surgical conditions, making it difficult to accurately identify anatomy and lesion sites, potentially leading to incorrect operations and reducing the success rate of surgery. Thus, laparoscopic surgical image restoration (LSIR) is critical to providing surgeons with clear and stable intraoperative views, significantly reducing surgery time, minimizing surgical errors, and improving operational efficiency.
Challenges: Unlike natural scenarios, the laparoscopic environment poses unique challenges for image restoration, characterized by diverse types and levels of image degradation, which seriously distort the surgical field of view and significantly hinder the distinction of anatomical structures [11, 23]. As illustrated in Fig. 1, Smoke originates from numerous fine tissue fragments and fat vaporization when surgeons use energy devices to cauterize diseased areas and surrounding tissues. Fog is mainly caused by condensation of water vapor when the temperature difference between the inside and outside of the abdominal cavity leads to mist forming on the lens. Compared to smoke noise, the distribution of surgical fog is broader and more uniform. For Splashing, the use of instruments such as electrocautery and ultrasonic scalpels during surgery results in the splashing of blood, tissue fluid, and other substances onto the lens, causing contamination and increasing the difficulty of restoration. Moreover, the restoration process requires cleaning anatomical details, e.g., blood vessels and tissue structures, to assist doctors in making precise surgical decisions. When complex degradations are superimposed, critical clues obscured by noisy substances may blend into the background, further increasing degradation complexity. To this end, LSIR presents unique complexities and challenges, highlighting the urgency of developing dedicated benchmark datasets for surgical restoration.
Figure 1: Illustration of surgical image restoration in laparoscopic surgery, including laparoscopic operating scenes, multi-type degradation causes and images, and algorithms for surgical restoration.
Datasets: Existing LSIR datasets usually have limitations in terms of authenticity and diversity. Several surgical restoration datasets [25, 4, 30] are generated through manual synthesis, resulting in discrepancies with real-world data and restricting the applicability of algorithms in actual scenarios. Other datasets derived from real surgical scenes [28, 21, 31] mainly focus on a single type of degradation, such as desmoking. To comprehensively investigate degradation challenges in real-world laparoscopic surgery, we construct the first multi-source LSIR dataset, SurgClean, aimed at effectively assisting surgeons in reducing visual noise in the operating field of view. SurgClean comprises 1,020 laparoscopic images under various degradation conditions selected from real surgical videos of 134 patients, accompanied by adjacent clean frames as paired reference labels. Driven by clinical needs, our dataset introduces three challenging LSIR tasks: desmoking, defogging, and desplashing, and further subdivides the degradation levels according to difficulty and characteristics. This dataset enables researchers to systematically explore and evaluate various image restoration methods under real laparoscopic conditions, providing opportunities for broader research and algorithm development.
Benchmarks: We establish a comprehensive LSIR benchmark to address the inherent data, properties, and computational challenges based on SurgClean. For desmoking, defogging, and desplashing, we evaluate 22 representative image restoration methods, including general and task-specific frameworks. Extensive experimental results reveal that: 1) Although existing image restoration techniques perform effectively in natural scenes, their accuracy and efficiency are insufficient to meet clinical requirements when dealing with complex multi-factor degradation. 2) While current methods show reasonable performance for samples with lower-level degradation, they fail to produce noticeable improvements for severe degradation at higher levels. These findings highlight the significant potential and necessity for developing advanced algorithms specifically designed for LSIR. By contributing to this benchmark community, we aim to guide future research prospects and methodological innovations, ultimately enhancing the efficiency of minimally invasive surgery and reducing surgical risks.
The main contributions of this study are three-fold: (1) We conducted a thorough exploration of various types of degradation in real-world surgical scenarios and introduced SurgClean, the first multi-source LSIR dataset in laparoscopic surgery. (2) We established a comprehensive LSIR benchmark based on the SurgClean dataset, achieving a performance evaluation of multiple types and levels of image restoration methods. (3) We revisited various degradation characteristics specific to surgical conditions and performed a manifold discrepancy analysis with degradation in natural scenes, intending to explore potential optimization strategies for surgical degradation.
2 Related Works
Figure 2: Illustration of degraded images and corresponding unaligned reference labels in SurgClean. The red arrows indicate noticeable ghosting artifacts in the overlapping regions.
Surgical Restoration Datasets. Various surgical image restoration (SIR) datasets [25, 4, 30, 28, 21, 33] have been developed to support the training and evaluation of restoration models in surgical scenarios. Among them, datasets such as [25, 4, 3, 30] are generated via manual synthesis during post-processing, leading to notable discrepancies from real-world surgical degradations and limiting their applicability to clinical practice. For instance, Chen et al. [4] proposed the SmokeGCN dataset, which applies Gaussian filters to generate various smoke effects and overlay them on clean images. While synthetic datasets have facilitated early research efforts in SIR, they struggle to adapt to the complexity and variability of clinical environments, thereby limiting the generalization ability of models. In contrast, datasets like Cyclic-DesmokeGAN [28], Desmoke-LAP [21], DesmokeData [33], and LSVD [31] are derived from real endoscopic surgeries and focus on addressing desmoking task. DeSmoke-LAP [21] focuses on desmoking in hysterectomy videos using an unpaired image-to-image translation approach. LSVD [31] comprises various smoke scenes from real laparoscopic surgery videos and adopts pre-smoke frames as supervision references. Although these datasets provide valuable resources for specific restoration tasks, they mainly focus on isolated scenarios and lack diversity in degradation types. To bridge this gap, we present a new comprehensive dataset to systematically explore multiple real-world surgical degradation challenges.
Image Restoration Approaches. General image restoration methods can be categorized into three categories: Convolutional Neural Network (CNN)-based, Transformer-based, and Mamba-based architectures. CNN-based methodologies like ConvIR [9] and FocalNet [8] demonstrate robust performance across diverse restoration tasks by utilizing multi-scale feature extraction capabilities. Transformer-based techniques such as Restormer [36], X-Restormer [5], AST [40], RAMiT [7], and Histoformer [27] improves restoration quality through attention mechanisms that enhance their modeling capacity. Recent advancements in state space models, particularly with the emergence of the Mamba architecture, have propelled Mamba-based modeling into prominence across various vision tasks including image restoration, with notable implementations like MambaIR [12] and MambaIRv2 [13]. As for the defogging task: RIDCP [32] proposes an enhanced haze imaging model integrated with codebook-based methodology to achieve robust image dehazing. C2PNet [38] implements contrastive learning to strengthen discriminative capability. DEA [6] focuses on designing a more effective network architecture to extract robust features. CORUN [10] develops a collaborative unfolding network, while CoA [20] employs a compress-then-adapt strategy to bridge synthetic-to-real domain gaps. For the desmoking task: DCP[14] serves as a classical approach that offers a simple yet effective solution for smoke removal. MS-CNN [29] integrates Laplacian pyramid decomposition with CNN architecture to perform desmoking in medical imaging scenarios. IIT-EDC [24] utilizes a lightweight GAN framework for real-time processing, while Desmoke-LAP[21] builds upon CycleGAN for medical smoke elimination. SelfSVD [31] implements self-supervision by utilizing clean frames from smoke-contaminated videos to guide the desmoking process.
3 SurgClean Dataset
Stemming from clinical requirements, our SurgClean dataset provides three challenging laparoscopic surgical image restoration (LSIR) tasks in real-world clinical scenarios, i.e., desmoking, defogging, and desplashing. We provide a detailed overview of SurgClean in terms of data collection, data annotation, and statistics analysis.
3.1 Data Collection
The degradation of surgical scenes caused by contaminants such as tissue fragments and smoke during laparoscopy occurs according to the specific surgical procedure and has a certain degree of randomness. Therefore, we focus only on restoring noise images that appear at specific time points rather than video-level restoration to improve clinical application efficiency. Concretely, we collect laparoscopic surgical videos of 134 patients undergoing hepatectomy and cholecystectomy surgeries from partner hospitals to construct the SurgClean dataset. Four intern surgeons were invited to pick out the frames from each video where they identified the surgical field as obscured by visual interference, including smoke, fog, and splash contamination on the lens. Meanwhile, two professional surgeons review and filter all interfering image frames to ensure the reliability of the samples. Ultimately, we gather a total of 1,020 noisy images with a resolution of 1280$\times$720, including 600 images for desmoking, 320 images for defogging, and 100 images for the desplashing task. The data in each sub-task is divided into training and test sets in an 8:2 ratio. Fig. 3 illustrates the data distribution across the three types of LSIR in SurgClean. The proportion of samples taken up by different sub-tasks also reflects the real-world occurrence rates of interference events during hepatectomy and cholecystectomy procedures. For instance, the probability of interference due to splashing is the lowest among the three types. Fig. 3 also exhibits various complex examples of visual interference encountered in laparoscopic surgery.
Figure 3: Illustration of the proposed SurgClean dataset. We exhibit data distribution under different degradation scenarios and further provide fine-grained divisions according to interference levels and attributes. Typical samples are displayed on both sides, zoom in for details.
3.2 Data Annotations
Since all samples in SurgClean are derived from real-world laparoscopic surgical videos, it is difficult to obtain perfectly aligned ground truth frames. To provide relatively clear paired reference labels for deep learning-based model supervision, we adopt the PS-frame scheme in line with [31], i.e., using the last clean frame preceding the noisy frame as the paired reference. As shown in Fig. 2, this form of label pair may be affected by the movement of the laparoscopic lens, resulting in slight deviations in pixel distribution. This also poses a key challenge for model supervision in this benchmark. To reduce the impact of misaligned references on the restoration model during training, we uniformly employ the optical flow alignment method [37, 26] to estimate the optical flow between the input image and the unaligned reference for relative displacement. Then, we warp the reference using the estimated optical flow to transform the unaligned images into aligned pairs. This strategy allows us to train and evaluate models as same as general image restoration tasks for fairness and effectiveness.
3.3 Data Statistics and Analysis
Table 1: Degradation level description.
Level Descriptions
Level 1
Mild degradation: Less than 1/3 of the entire field of view is filled
with smoke or fog, and the anatomical structures in the operating area
are visible and do not affect judgment.
Level 2
Moderate degradation: 1/3-2/3 of the entire field of view is filled with
smoke or fog, and the anatomical structures in the operating area are
still visible but less clearly than in Level 1, without affecting judgment.
Level 3
Severe degradation: More than 2/3 of the entire field of view is filled
with smoke or fog, or the operating area
is degraded, which may affect judgment.
Level 4
Complete degradation: More than 2/3 of the entire field of view is filled
with smoke or fog, and the operating area is degraded, severely affecting
judgment. Surgery must be paused to wait for the degradation to clear.
To facilitate more fine-grained research, we further divided the dataset into multiple subsets based on different degradation levels and attributes. Specifically, for desmoking and defogging tasks, surgeons divide corresponding samples into four levels (from easy to difficult, i.e., Level 1-4) based on the area and degree of smoke or fog obstruction and the degree of interference with the operating zone. For desplashing, the samples are grouped into four subsets (i.e., $T_{blood}$, $T_{fat}$, $T_{bile}$, and $T_{fluid}$) based on the type of substances splashed onto the lens, including blood, fat, bile, and tissue fluid. We clearly elaborate on the definition and classification of different levels of degradation in Table 1. In addition, Fig. 3 exhibits typical examples of different difficulty levels and each type of splash interference. More sample exhibitions and information statistics of our SurgClean dataset can be found in Appendix. B.
Table 2: Comparison between SurgClean and existing real-world surgical image restoration datasets.‘OA’ stands for whether the data are publicly available.
Dataset OA Paired Label Resolution Desmoke Defog Desplash Scale Source No. Source
Cyclic-DesmokeGAN [28] 240$\times$320 1,400 1 Cholec80
Desmoke-LAP [21] 720$\times$540 3,000 1 Hysterectomy
DesmokeData [33] 700$\times$350 961 1 da Vinci Si videos
SurgClean (Ours) 1280$\times$720 1,020 4 Cholecystectomy, Hepatectomy, Pancreatectomy, Splenectomy
Furthermore, we summarize and perform comparisons with other real-world LSIR datasets. As illustrated in Table 2, existing LSIR datasets focus only on the desmoking task, while SurgClean comprehensively embraces multiple types of image degradation in actual surgical procedures. In addition, datasets with paired reference labels provide more significant support for model fine-tuning. Compared to other datasets with a single data source, our data source comes from various types of laparoscopic surgery, e.g., hepatectomy, cholecystectomy, and splenectomy, to accommodate data diversity. SurgClean comprises a variety of interference scenarios in laparoscopic surgery to promote comprehensive evaluation and advancement in surgical image restoration tasks. Additionally, the presence of paired reference labels facilitates the training and evaluation of deep learning-based restoration models. We conduct cross-validation experiments with DesmokeData [33] to validate the effectiveness and challenging nature of our dataset (refer to Appendix. E). Meanwhile, we also perform a visual comparison of degraded images in our dataset with other LSIR datasets in Appendix. C. Our dataset has passed the ethics statement and can be publicly available at https://github.com/PJLallen/Surgical-Image-Restoration.
4 Benchmarking and Results
We provide an overview of benchmark construction and evaluation for surgical image restoration (SIR) based on SurgClean dataset. The benchmark includes implementation details, evaluation metrics, training setting, and performance evaluation and analysis of various representative image restoration models, including generic and task-specific image restoration models. All pre-trained weights and evaluation codes of our benchmark are available at https://github.com/PJLallen/Surgical-Image-Restoration.
4.1 Experimental Setup
Implementation Details. All methods are implemented with PyTorch [22] and trained on two NVIDIA RTX 4090 GPUs with the Adam optimizer. During training, the input image is randomly cropped into 128$\times$128 patches with a batch size of 2 and a total of 200,000 iterations. In addition, we use random crop and horizontal flips as data augmentation. The initial learning rate of all methods is set to default value, and the learning rate is decreased by half every 100k iterations throughout the training process. Additionally, all models are trained on the training set of each sub-scenario and then individually inferred on the corresponding test set on SurgClean. For desmoking, 480 images are employed for training and 120 for testing. For defogging, 256 images are used for training and 64 for testing. For desplashing, the subset divides the training and test samples in an 8:2 ratio.
Evaluation Metrics. In line with general evaluation metrics for image restoration tasks, this benchmark evaluates model performance using Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) metrics. Referring to [31], we also adopt non-reference metrics, i.e., Naturalness Image Quality Evaluator (NIQE) and Perceptual Index (PI), to evaluate the restoration results of three sub-degradation scenarios.
Training Setting. As mentioned in Sec. 3.2, our dataset provides unaligned paired reference labels to facilitate the training process of restoration models. To mitigate the impact of unaligned references, all comparison models use pre-trained PWC-Net [26] to estimate the optical flow of unaligned reference $\mathbf{UR}$ so that it aligns with the predicted restored image $\mathbf{P}$:
$\mathbf{F}_{{UR}\rightarrow P}=\mathcal{O}(\mathbf{UR},\mathbf{P}),$ (1)
where $\mathcal{O}$ denotes the pre-trained optical flow network. Given the estimated optical flow $\mathbf{F}_{{UR}\rightarrow P}$, we apply the warping operation $\mathcal{W}$ to the $\mathbf{UR}$ image, transforming the unaligned label into aligned $\mathbf{UR}_{warp}$:
$\mathbf{UR}_{warp}=\mathcal{W}(\mathbf{UR},\mathbf{F}_{UR\rightarrow P}).$ (2)
Finally, $\mathbf{UR}_{warp}$ is spatially aligned with the restored image $\mathbf{P}$. The reconstruction loss $\mathcal{L}_{rec}$ of the restoration model can be expressed as
$\mathcal{L}_{rec}=\sum_{i=1}^{N}||\mathbf{M}_{i}\odot(\mathbf{UR}_{warp,i}-\mathbf{P}_{i}||_{1}.$ (3)
where $\odot$ is the pixel-wise multiplication, and Mi masks areas with imperfect optical flow to prevent trivial solutions. This allows the model to be trained and evaluated in the same way as the image restoration task with aligned GT. We apply this warping-based training and evaluation strategy to all models in our benchmark to ensure fairness and effectiveness.
Table 3: Quantitative comparison for surgical image restoration with representative generic restoration models on the SurgClean test sets with PSNR $\uparrow$, SSIM $\uparrow$, LPIPS $\downarrow$, NIQE $\downarrow$ and PI $\downarrow$ metrics.
Methods Pubs Desmoking Defogging Desplashing Params
PSNR SSIM LPIPS NIQE PI PSNR SSIM LPIPS NIQE PI PSNR SSIM LPIPS NIQE PI
Restormer [36] CVPR’22 18.470 0.634 0.368 5.373 11.854 18.154 0.619 0.470 6.757 30.084 21.732 0.708 0.302 5.111 25.807 26.13M
FocalNet [8] ICCV’23 18.347 0.655 0.397 5.631 36.676 18.250 0.626 0.474 7.066 41.550 21.740 0.722 0.286 5.557 35.590 3.74M
ConvIR [9] TPAMI’24 18.464 0.649 0.399 5.807 42.004 18.022 0.617 0.471 6.782 33.173 21.745 0.720 0.309 5.476 37.437 14.83M
Fourmer [39] ICML’24 18.208 0.642 0.408 5.418 28.154 17.712 0.610 0.484 6.530 32.162 21.597 0.714 0.300 5.327 30.685 0.68M
MambaIR [12] ECCV’24 18.357 0.649 0.418 6.310 43.049 17.338 0.625 0.524 7.801 48.292 21.908 0.727 0.310 6.411 46.272 4.31M
Histoformer [27] ECCV’24 16.687 0.376 0.640 6.939 13.202 16.183 0.348 0.632 7.393 11.799 20.348 0.637 0.320 5.408 2.348 29.92M
RAMiT [7] CVPR’24 18.348 0.647 0.386 5.841 29.850 18.162 0.622 0.483 7.070 40.173 21.662 0.716 0.305 5.424 34.413 0.30M
AMIR [35] MICCAI’24 18.097 0.637 0.409 6.332 17.141 18.191 0.604 0.521 7.266 48.706 21.759 0.701 0.285 4.962 11.363 23.54M
AST [40] CVPR’24 18.630 0.617 0.405 5.317 9.088 16.572 0.590 0.422 6.461 25.395 21.865 0.720 0.302 4.977 25.109 19.92M
X-Restormer [5] ECCV’24 17.570 0.647 0.398 5.545 37.243 18.148 0.615 0.471 6.972 26.732 21.882 0.727 0.307 5.575 37.966 42.52M
SFHformer [15] ECCV’24 17.965 0.637 0.405 5.688 9.833 18.035 0.613 0.483 7.168 15.285 21.913 0.727 0.296 5.438 34.271 7.67M
MambaIRv2 [13] CVPR’25 18.572 0.638 0.422 6.046 44.250 17.942 0.625 0.519 7.878 53.505 21.995 0.728 0.315 5.761 44.683 0.77M
4.2 Benchmark Approaches
We conduct a comprehensive evaluation of 22 representative image restoration methods on SurgClean, including 12 generic image restoration models and 10 task-specific image restoration models. Specifically, generic restoration frameworks are built on different infrastructures, including CNN-based ConvIR [9] and FocalNet [8], transformer-based Restormer [36], X-Restormer [5], AST [40], RAMiT [7], and Histoformer [27], state space model-based MambaIR [12] and MambaIRv2 [13], and Fourier-based methods Fourmer [39] and SFHformer [15].
Besides, we employ task-specific frameworks to validate the effectiveness in single degradation scenarios, including MS-CNN [29], Desmoke-LAP [21], IIT-EDC [24], DCP [14], and SelfSVD [31] for desmoking, and C2PNet [38], RIDCP [32], CORUN [10], DEA [6], and CoA [20] for defogging.
Figure 4: Quantitative comparison for surgical image restoration with representative task-specific restoration models on the SurgClean test sets. The upper and lower panels comparatively demonstrate defogging and desmoking, respectively.
Figure 5: Qualitative comparison for surgical image restoration on the SurgClean dataset. The methods selected for visual comparison were chosen from the top-five performing approaches in Table 3, based on their quantitative benchmark metrics.
4.3 Benchmark Results
Quantitative Analysis. We conduct a comprehensive quantitative comparison of various representative generic and task-specific image restoration methods on the SurgClean dataset for desmoking, defogging, and desplashing tasks. As shown in Table 3, existing generic restoration methods still face unique challenges when dealing with complex degradation issues in real-world surgical environments. Specifically, AST [40] achieves relatively superior PSNR, NIQE, and PI scores, i.e., 18.630, 5.317, and 9.088, respectively for desmoking, but these results remain insufficient for high-quality reconstruction of smoke degradation during surgery. FocalNet [8] achieves higher PSNR and SSIM values in the defogging task, but lower NIQE and PI values indicate that restored images still fall short of clean images in terms of authenticity and naturalness. For desplashing, MambaIRv2 [13] and SFHformer [15] perform better in terms of PSNR and SSIM metrics, but the results require further improvement to meet the standards for clinical applications. In comparison, the RAMiT framework [7] has only 0.3M model parameters, which provides certain advantages when in deployment on actual surgical equipment. More fine-grained quantitative results can be found in Appendix. D.1.
Fig. 4 illustrates the quantitative performance of task-specific restoration models for desmoking and defogging on the SurgClean test sets. We can see that approaches specialized for desmoking do not yield better restoration performance compared to generic image restoration models, such as MS-CNN [29] with a score of 18.48 compared to 18.63 achieved by AST [40]. For the defogging task, task-specific models also perform slightly worse than generic restoration models, highlighting the uniqueness and complexity of fog degradation in laparoscopic scenes. Overall, although various restoration techniques have developed rapidly in recent years, the specificity of degradation in multiple types of surgical scenarios, e.g., structural diversity and local damage, hinders the generalization of existing restoration models to real surgical environments.
Qualitative Analysis. We provide visual analysis of several restoration methods with favorable quantitative results in typical desmoking, defogging, and desplashing scenarios. As shown on the left side of Fig. 5, most methods exhibit limited effectiveness in removing dense smoke within localized operating areas, and residual smoke still obscures anatomical structures. FocalNet [8] and AST [40] perform better on tissue visibility, but still struggle to recover fine anatomical details. For defogging, restoration models fail to completely suppress fog to provide a clear view, and structural distortion remains in the operating area. Similarly, most approaches are ineffective in eliminating occlusions caused by local splashes. In general, all benchmarking methods show limited performance when addressing various types of degradation in surgical scenarios, highlighting the uniqueness and persistent LSIR challenges. More visualization results can be found in Appendix. D.2.
5 Beyond Pixel-Level Recovery
In this section, we attempt to look beyond the “pixel-level recovery” perspective from two aspects, including scene analysis between surgical and natural environments and surgical scene parsing.
5.1 Scene Analysis between Surgical and Natural Environments
Although Section 4 explores a direct adaptation strategy by leveraging existing natural image restoration models for surgical scenarios, extensive experiments reveal fundamental incompatibilities. In this part, we try to explore its fundamental cause. As illustrated in Fig. 6 (left), the t-SNE visualization demonstrates significant distributional divergence between natural and surgical scenes (natural haze [17], natural dense haze [1, 2]), indicating inherent domain gaps at the statistical level. Furthermore, we utilized the depth estimation model [34] to perform depth-related scattering degradation analysis (Fig. 6 (right)), highlighting critical imaging distinctions, i.e., laparoscopic fog exhibits localized depth variations with abrupt transitions, whereas natural scenes present gradual depth progressions across expansive environments. These compounded discrepancies, which span both statistical feature spaces and physical imaging degradation characteristics, collectively explain why conventional natural image restoration paradigms struggle in surgical domains.
Figure 6: Scene analysis between surgical and natural environments. t-SNE feature visualization (Left) and depth estimation examples (Middle/Right).
Figure 7: Surgical scene parsing: SAM [16] (Top) and MedSAM [19] (Bottom).
5.2 Surgical Scene Parsing
While visual data quality remains essential from clinical perspectives, the integration of intelligent assistive cues (e.g., scene understanding) proves critical for advancing precision medicine in surgical practice. To systematically evaluate how image restoration quality impacts downstream tasks, we employed the SAM [16] for scene segmentation and its medical variant MedSAM [19] for instrument localization across surgical scenarios. As demonstrated in Fig. 7, although different restoration algorithms have brought improvements in visual quality, their performance in segmentation tasks does not show stable and consistent performance improvement. Especially in surgical scenes with complex structures and severe occlusion, it is still difficult to achieve accurate segmentation of tissue structures and surgical instruments. Therefore, achieving the optimal equilibrium between visual fidelity and downstream task-specific performance emerges as a critical challenge for maximizing the clinical utility of surgical image restoration models. This fundamental trade-off underscores the necessity of semantic-aware restoration frameworks, charting a promising research direction.
6 Discussion and Conclusion
Discussion. The core objective of this study is to construct a multi-type laparoscopic surgical image restoration (LSIR) benchmark dataset for complex surgical scenarios. By establishing a standardized evaluation platform, we aim to drive the development of restoration algorithms to meet clinical needs. Here, we present a deeper analysis and insights into SurgClean benchmark:
1) Suboptimal benchmark performance. Both quantitative and qualitative analyses indicate that existing restoration methods are inadequate for handling the highly complex and variable degradations in surgical environments. Compared to natural image restoration, LSIR poses unique challenges, e.g., artifact blending with the background, localized occlusions, and the entanglement of diseased tissues with degraded areas, which significantly increase the complexity of surgical scenes. We attempted to explore the differences between surgical and natural images from the perspectives of structural depth and semantic understanding to provide insights for future research.
2) Selection of degradation types and data scale. We selected three degradation scenarios (i.e., smoke, fog, and splashing) from 134 real surgical cases that experienced surgeons identified as directly affecting surgical efficiency. The distribution of samples across different degradation types reflects their relative frequency in the full dataset. We plan to expand the dataset with more real-world samples to facilitate further research in the LSIR community.
3) Simultaneous multi-type degradation. While it is theoretically possible for multiple degradation types to occur simultaneously in the same scenario, such combined degradations were rarely observed in our data collection process. This may be due to differing phases and conditions under which the three types of degradation occur. Thus, the first multi-type surgical restoration benchmark only focuses on samples with a single degradation type in a single scene. Future work will explore simultaneous degradations to enhance the LSIR diversity and challenge.
4) Clinical significance. The surgical image restoration task provides surgeons with a clear surgical field of view, thereby significantly reducing operation time, minimizing surgical errors, and improving procedural efficiency. By eliminating unnecessary visual obstructions, surgeons can maintain focus during operations. Beyond enhancing image quality, LSIR also involves reconstructing complex anatomical details, which is critical for precise surgical navigation and shortening postoperative recovery times for patients.
5) Future research prospects. Given the unique challenges of LSIR, future research can generate large amounts of synthetic data by simulating the imaging principles of surgical degradation to further reduce training costs. Besides, distilling semantic knowledge from large foundation models could improve the identification and localization of degraded regions. Finally, developing lightweight networks with optimized computational efficiency ensures practical deployment in clinical surgery.
Conclusion. We propose a comprehensive real-world laparoscopic surgical image restoration (LSIR) benchmark to address clinical demands in mitigating various scene degradation distortions. We introduce SurgClean, the first multi-type LSIR dataset tailored for real surgical scenarios. SurgClean provides 1,020 images for desmoking, defogging, and desplashing, as well as fine-grained divisions based on degradation levels and categories. We establish a standardized evaluation protocol and conduct systematic benchmark testing. The results illustrate that existing restoration techniques still fall short of meeting clinical requirements for a clear visual field. Beyond pixel-level perspective, we explore the structural differences and unique characteristics between surgical and natural scene degradations, and analyze the broad influence of restoration quality on downstream surgical scene understanding tasks. Through the proposed SurgClean benchmark, we aim to foster the emergence of future research to advance surgical restoration technology and ultimately contribute to reliable intraoperative visual assistance to reduce surgical risks.
Appendix A Overview
Our supplementary material includes the following sections:
  • Section B: SurgClean Dataset. More sample exhibitions and information statistics about the proposed SurgClean dataset.
  • Section C: Visual Analysis with LSIR Datasets. Data visual comparison with other open-source laparoscopic surgical image restoration datasets.
  • Section D: More Benchmarking Results. The comparative results on fine-grained subsets of SurgClean and additional visualization results.
  • Section E: Cross-validation between Datasets. Cross-validate the performance of image restoration models on different laparoscopic surgical restoration datasets.
  • Section F: Different Training Strategies. The effect of ablating different pre-training weights and training strategies on restoration performance.
  • Section G: Limitations. Discussion of limitations in this work.
  • Section H: License of the Used Assets.
Appendix B SurgClean Dataset
B.1 Clarity for Label Alignment
As mentioned in Section 3.2, it is impossible to obtain perfectly aligned reference labels that match degraded images at the exact same moment in real-world laparoscopic surgical videos. Therefore, we extract frames from the video just before the scene becomes degraded to obtain an approximate paired image, which inevitably introduces misalignment. The misalignment in our study is the same as the issue in LSVD [31]. Therefore, the optical flow alignment method based on PWC-Net [26] is equally applicable to our benchmark. Fig. 8 exhibits the visualization of reference labels before and after alignment on SurgClean dataset. It can be observed that the optical flow-based warping method effectively converts misaligned images to alignments and masks out the error regions.
B.2 More Data Visualization
To facilitate a more comprehensive understanding, analysis, and interpretation of SurgClean, we present more fine-grained data visualization for Defogging, Desmoking and Desplashing tasks in Fig. 9, Fig. 10, and Fig. 11. As can be seen, the multi-type degradation samples in SurgClean are hierarchical, and the diverse degradation scenarios also reflect the great challenge of the real laparoscopic surgical image restoration task.
Figure 8: Visual comparison for reference label alignment on the SurgClean dataset.
Figure 9: More fine-grained data exhibitions for desmoking in SurgClean dataset.
Figure 10: More fine-grained data exhibitions for defogging in SurgClean dataset.
Figure 11: More fine-grained data exhibitions for desplashing in SurgClean dataset.
B.3 Basic Information of Annotated Patients
In Table 4, we present comprehensive demographic, clinical, and surgical information from 134 patients who underwent laparoscopic surgery at the department of hepatobiliary surgery in our partner hospital. The dataset consists of 78 males and 56 females with an average age of 52±19 years.
Surgical Procedures and Types. The dataset encompasses a diverse range of surgeries, categorized as gallbladder, bile duct, liver, pancreas, and spleen surgeries.
Among gallbladder surgeries, cholecystectomy for inflammation involved 8 males and 15 females, while cholecystectomy for gallstones, polyps, or adenomyomatosis involved 19 males and 17 females. The median duration for these procedures was 40 and 30 minutes, respectively.
For bile duct surgery, common bile duct exploration was performed on 2 males and 3 females, with a median duration of 100 minutes. Liver surgeries were more varied, including cyst fenestration (1 male, 2 females, 50 minutes), radical hepatectomy for hepatocellular carcinoma (29 males, 11 females, 170 minutes), hemangioma resection (2 males, 6 females, 150 minutes), radical cholecystectomy for gallbladder cancer (1 male, 200 minutes), and focal nodular hyperplasia resection (5 males, 120 minutes).
Pancreatic surgeries consisted of distal pancreatectomy (3 males, 1 female, 220 minutes) and pancreatoduodenectomy (5 males, 1 female, 420 minutes), while spleen surgeries involved splenectomy for portal hypertension (3 males, 160 minutes).
Surgical Outcomes. Among liver surgery patients, cirrhosis was a common comorbidity, with 14 patients diagnosed preoperatively or intraoperatively. Extensive liver resection, required for large tumors, was associated with a higher risk of postoperative liver failure. Nine patients underwent extensive resections, with a median surgical duration of 210.5 minutes and median blood loss of 220 ml. The median postoperative hospitalization duration was 8 days, with one patient experiencing liver failure that was resolved prior to discharge.
Table 4: Comprehensive information of annotated patients from SurgClean Dataset.
Procedure Category Procedure Type Male Female Median Surgery Duration (min)
Gallbladder Surgery Inflammation (Cholecystectomy) 8 15 40 (20)
Gallstones/Polyps/Adenomyomatosis (Cholecystectomy) 19 17 30 (15)
Bile Duct Surgery Common Bile Duct Exploration 2 3 100 (30)
Liver Surgery Cyst Fenestration 1 2 50 (20)
Radical Hepatectomy (Hepatocellular Carcinoma) 29 11 170 (60)
Hemangioma Resection 2 6 150 (25)
Radical Cholecystectomy (Gallbladder Cancer) 1 0 200
Focal Nodular Hyperplasia Resection 5 0 120 (30)
Pancreas Surgery Distal Pancreatectomy 3 1 220 (40)
Pancreatoduodenectomy 5 1 420 (70)
Spleen Surgery Splenectomy (Portal Hypertension) 3 0 160 (20)
Total - 78 56 -
Figure 12: Visual Comparison with LSIR Datasets.
Appendix C Visual Analysis with LSIR Datasets
We conducted a comprehensive visual analysis comparing the SurgClean dataset with existing LSIR benchmarks, specifically evaluating Desmoke-LAP [21] and DesmokeData [33]. (Note that Cyclic-DesmokeGAN [28] was excluded from this comparison due to its non-public availability.) As illustrated in Figure 12, critical distinctions emerge between these datasets. Both Desmoke-LAP and DesmokeData exhibit inherent limitations: they are generated from single surgical sources, which restricts their applicability exclusively to the desmoking task. Furthermore, Desmoke-LAP suffers from the additional constraint of containing unpaired training samples. At the same time, Desmoke-LAP and DesmokeData provide lower-resolution images that may hinder model performance in fine-grained surgical scenarios.
In contrast, SurgClean demonstrates significant advancements in dataset design. First, it incorporates meticulously paired data collected across four distinct surgical procedures, including Cholecystectomy, Hepatectomy, Pancreatectomy, and Splenectomy, ensuring broader clinical relevance. Second, our dataset supports three advanced vision restoration tasks (i.e., desmoking, defogging, and desplashing), surpassing the single-task focus of existing alternatives. Third, SurgClean provides relatively high-resolution imaging data that preserves critical details, enhancing key limitations of prior datasets.
Appendix D More Benchmarking Results
D.1 Fine-grained evaluation on sub-test sets of SurgClean
To further validate the performance of representative methods, we also report PSNR and SSIM values of sub-test sets for three restoration tasks on SurgClean in Table 5 and Table 6, i.e., Level 1–4 for desmoking and defogging, and bile, blood, fat, and tissue fluid for desplashing. The fine-grained evaluation highlights several important findings. First, in the desmoking and defogging tasks, as the degradation level increased, the restoration performance of evaluation models consistently declined, indicating that existing methods exhibit insufficient restoration robustness when the degradation level increases. For example, while FocalNet [8] achieves the best PSNR (21.2585) at Level-1 desmoking, its performance drops sharply to 16.2681 at Level-4. Second, for fine-grained desplashing, there are significant differences between different splash substances. MambaIR [12] and MambaIRv2 [13] show relatively stable restoration results across different sub-test sets. However, no single model dominates all levels and types of degradation scenarios. Overall, the fine-grained experimental results indicate that laparoscopic surgical image restoration still faces significant challenges under diverse and severe degradation conditions.
Table 5: Fine-grained comparison with generic restoration models for desmoking and defogging on sub-test sets of SurgClean dataset.
Tasks Methods Publications Level-1 Level-2 Level-3 Level-4
PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
Desmoking Restormer [36] CVPR’22 21.1306 0.6913 18.4065 0.6453 16.4002 0.5489 15.5004 0.6119
FocalNet [8] ICCV’23 21.2585 0.7008 18.3774 0.6601 16.5798 0.5425 16.2681 0.6295
ConvIR [9] TPAMI’24 21.1700 0.7012 18.4728 0.6561 16.3844 0.5557 15.1461 0.6367
Fourmer [39] ICML’24 20.3133 0.6980 18.2348 0.6552 16.5035 0.5552 15.8499 0.6253
MambaIR [12] ECCV’24 20.5822 0.7059 18.3583 0.6608 16.7229 0.5578 15.9875 0.6364
Histoformer [27] ECCV’24 19.3702 0.4471 16.0145 0.4303 14.0606 0.3719 13.0636 0.3938
RAMiT [7] CVPR’24 21.2434 0.7057 18.2444 0.6590 16.1250 0.5555 15.5262 0.6305
AMIR [35] MICCAI’24 20.9472 0.6987 18.1020 0.6504 16.0858 0.5447 14.9289 0.6113
AST [40] CVPR’24 21.1211 0.6776 18.6088 0.6281 16.7118 0.5289 15.8765 0.5850
X-Restormer [5] ECCV’24 21.0098 0.7119 17.4058 0.6601 15.2322 0.5550 13.4543 0.6074
SFHformer [15] ECCV’24 21.2009 0.6994 17.8233 0.6495 15.6640 0.5463 14.4804 0.6111
MambaIRv2 [13] CVPR’25 20.8124 0.6975 18.4392 0.6448 16.8815 0.5531 16.0999 0.6197
Defogging Restormer [36] CVPR’22 19.5814 0.6535 18.6327 0.6247 17.7253 0.6046 16.4570 0.6125
FocalNet [8] ICCV’23 19.5731 0.6415 18.6499 0.6187 17.8005 0.6049 16.4591 0.6043
ConvIR [9] TPAMI’24 19.2928 0.6472 18.5396 0.6204 17.8003 0.6087 16.3916 0.6054
Fourmer [39] ICML’24 18.8020 0.6315 18.2619 0.6117 17.5112 0.5986 16.3083 0.6142
MambaIR [12] ECCV’24 17.3317 0.6281 17.7685 0.6159 17.2433 0.6274 16.5664 0.6344
Histoformer [27] ECCV’24 18.0907 0.2963 17.0857 0.2569 16.6102 0.2668 15.2006 0.2517
RAMiT [7] CVPR’24 19.7547 0.6547 18.5891 0.6254 17.8099 0.6151 16.3768 0.6054
AMIR [35] MICCAI’24 19.7000 0.6412 18.5595 0.5990 17.8647 0.6049 16.3827 0.5823
AST [40] CVPR’24 17.5711 0.5887 16.6786 0.5636 16.4697 0.6076 15.3710 0.5840
X-Restormer [5] ECCV’24 19.5498 0.6528 18.7151 0.6179 17.8697 0.6079 16.7165 0.5986
SFHformer [15] ECCV’24 19.5287 0.6390 18.6425 0.6177 17.6314 0.5977 16.4193 0.5898
MambaIRv2 [13] CVPR’25 18.8396 0.6606 18.3663 0.6244 17.7639 0.6172 16.8245 0.6242
Table 6: Fine-grained comparison with generic restoration models for desplashing on sub-test sets of SurgClean dataset.
Tasks Methods Publications Bile Blood Fat Tissue Fluid
PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
Desplashing Restormer [36] CVPR’22 20.3297 0.7119 21.2885 0.6857 23.7017 0.7457 21.6298 0.7099
FocalNet [8] ICCV’23 20.2936 0.7145 21.2441 0.6838 23.7671 0.7641 21.5938 0.7143
ConvIR [9] TPAMI’24 20.2828 0.7112 21.1375 0.6849 23.7890 0.7653 21.6652 0.7202
Fourmer [39] ICML’24 20.0813 0.7056 21.1763 0.6705 23.6405 0.7608 21.5036 0.7070
MambaIR [12] ECCV’24 20.4183 0.7149 21.3762 0.6909 23.9665 0.7689 21.7875 0.7260
Histoformer [27] ECCV’24 19.2512 0.6419 19.5642 0.6003 22.3528 0.6687 20.1103 0.6298
RAMiT [7] CVPR’24 20.2766 0.7041 21.1948 0.6795 23.5274 0.7604 21.5724 0.7139
AMIR [35] MICCAI’24 20.2477 0.6968 21.2682 0.6652 23.7702 0.7411 21.6691 0.6940
AST [40] CVPR’24 20.2490 0.6984 21.4021 0.6829 23.8603 0.7588 21.6277 0.7197
X-Restormer [5] ECCV’24 20.2838 0.7095 21.2440 0.6867 23.8657 0.7666 21.7215 0.7216
SFHformer [15] ECCV’24 20.4714 0.7088 21.3538 0.6854 23.8002 0.7680 21.6545 0.7213
MambaIRv2 [13] CVPR’25 20.3860 0.7059 21.4479 0.6888 23.8165 0.7648 21.8318 0.7264
D.2 Additional Visual Results
We provide more qualitative results of representative image restoration models with superior quantitative results for desmoking, defogging, and desplashing on the SurgClean test set, including ConvIR [9], FocalNet [8], AST [40], Restormer [36], RAMiT [7], AMIR [35], MambaIR [12], SFHformer [15]. As shown in Fig. 13 and Fig. 14, the restoration models are relatively effective at lower degradation levels. When the smoke and fog levels reach 3 or even 4, the restoration visualization effect is no longer noticeable. It indicates that the noisy samples in the SurgClean dataset pose significant challenges for surgical image restoration. As shown in Fig. 15, for the brand-new desplashing task, all evaluated models did not show obvious responses to degraded regions. It reflects that desplashing is more challenging than the other two restoration tasks, which drives us to explore more specific algorithms in follow-up research to solve multi-type splashing problems in surgical restoration domain.
Figure 13: Additional visualizations of representative image restoration models for desmoking on SurgClean test set.
Figure 14: Additional visualizations of representative image restoration models for defogging on SurgClean test set.
Figure 15: Additional visualizations of representative image restoration models for desplashing on SurgClean test set.
Appendix E Cross-validation between LSIR Datasets
To further evaluate the differences between our dataset and other open-source surgical image restoration datasets, we also invite publicly available DesmokeData dataset  [33] with paired labels for cross-validation on the desmoking task. DesmokeData contains a total of 961 degraded images along with corresponding paired labels. We randomly divided 761 images for training and 200 images for testing. As shown in Table 7, we selected representative restoration models and conducted four sets of cross-validation experiments (Exp 1-4). Comparing the results of Exp 1 and Exp 2, we can see that almost all methods perform better on DesmokeData than on SurgClean. This indicates that the SurgClean dataset is significant higher than DesmokeData in terms of degradation complexity and restoration difficulty. The reason probably lies in the presence of multi-level dense smoke occlusion and more dynamic interference factors in SurgClean. When the restoration models are trained on DesmokeData and then tested on SurgClean, most methods showed a significant decline in performance compared to those trained on in-domain data. In contrast, in Exp 4, the models trained on SurgClean perform more robustly when tested on DesmokeData. This experiment demonstrates that models trained on SurgClean have stronger generalization performance, further highlighting the potential of our dataset in advancing the development of robust models.
Table 7: Cross validation for desmoking between DesmokeData [33] and SurgClean datasets.
Tasks Methods Training Set Test Set PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ NIQE $\downarrow$ PI $\downarrow$
Exp 1 ConvIR [9] DesmokeData DesmokeData 20.095 0.760 0.182 5.501 25.231
FocalNet [8] 19.185 0.744 0.193 5.462 12.165
Restormer [36] 22.027 0.782 0.148 5.399 31.882
MambaIR [12] 23.181 0.767 0.132 5.270 38.320
Histoformer [27] 23.594 0.772 0.129 4.832 21.106
RAMiT [7] 23.609 0.786 0.131 5.011 30.975
AMIR [35] 23.780 0.781 0.132 4.683 30.894
Exp 2 ConvIR [9] SurgClean SurgClean 18.464 0.649 0.399 5.807 42.004
FocalNet [8] 18.347 0.655 0.397 5.631 36.676
Restormer [36] 18.470 0.634 0.368 5.373 11.854
MambaIR [12] 18.357 0.649 0.418 6.310 43.049
Histoformer [27] 16.687 0.376 0.640 6.939 13.202
RAMiT [7] 18.348 0.647 0.386 5.841 29.850
AMIR [35] 18.097 0.637 0.409 6.332 17.141
Exp 3 ConvIR [9] DesmokeData SurgClean 15.920 0.525 0.435 5.204 19.885
FocalNet [8] 16.700 0.617 0.376 4.910 14.540
Restormer [36] 14.713 0.531 0.423 5.139 18.721
MambaIR [12] 15.782 0.580 0.400 5.644 35.077
Histoformer [27] 15.824 0.572 0.415 5.097 8.165
RAMiT [7] 15.306 0.560 0.396 5.071 19.372
AMIR [35] 15.167 0.557 0.411 5.184 15.842
Exp 4 ConvIR [9] SurgClean DesmokeData 24.099 0.853 0.158 4.112 10.235
FocalNet [8] 24.577 0.881 0.158 3.821 9.801
Restormer [36] 23.315 0.846 0.159 4.523 11.264
MambaIR [12] 22.405 0.768 0.231 5.487 14.981
Histoformer [27] 24.034 0.654 0.313 6.091 16.932
RAMiT [7] 23.503 0.823 0.156 4.328 10.706
AMIR [35] 22.464 0.865 0.186 5.073 12.823
Appendix F Performance of Different Training Strategies
While the image restoration task has been extensively studied in natural scenes, its application to surgical scenarios presents unique challenges due to the distinct characteristics of laparoscopic imaging. To systematically evaluate cross-domain generalization capabilities, we design a two-stage experimental scheme to validate the effect of natural and surgical scene data on the performance of restoration models in surgical defogging. In the first stage, we implement 12 representative restoration algorithms pre-trained on Haze4K [18], a synthetic dataset optimized for haze removal in natural scenes, and directly deploy them on the SurgClean sub-test set for defogging without any adaptation. As quantified in Table 8, these models exhibit substantial performance degradation, revealing significant domain discrepancies between natural and surgical fog patterns.
To investigate domain adaptation potential, the second stage involves fine-tuning the pre-trained model using the defogging training set of SurgClean while maintaining their architectural configurations. The result confirms that surgical-specific data are indispensable for bridging domain gaps through targeted retraining. These findings further emphasize the necessity of developing specialized benchmarks for laparoscopic scene restoration, as generic computer vision solutions prove insufficient for meeting the complex demands of clinical surgery.
Table 8: Comparing different training strategies for defogging on the SurgClean test sets.
Methods Trained on Natural Fog Finetuned on Surgical Fog
[…]
Restormer [36] 17.421 0.601 0.462 5.950 27.977 17.993 0.634 0.420 6.164 16.690
FocalNet [8] 15.842 0.572 0.473 6.019 32.298 18.125 0.627 0.476 7.230 40.204
ConvIR [9] 15.297 0.545 0.481 5.877 31.047 17.531 0.606 0.467 6.824 31.664
Fourmer [39] 17.252 0.603 0.452 5.802 30.844 17.609 0.620 0.471 6.656 33.994
MambaIR [12] 16.700 0.594 0.467 6.219 32.232 18.013 0.627 0.503 7.980 52.532
Histoformer [27] 16.849 0.593 0.461 5.783 22.822 15.948 0.540 0.470 5.503 4.995
RAMiT [7] 16.495 0.588 0.466 6.046 29.304 18.126 0.624 0.481 7.305 33.935
AMIR [35] 16.986 0.597 0.463 6.067 29.206 17.922 0.625 0.466 6.681 41.265
AST [40] 16.703 0.577 0.411 5.648 15.949 18.344 0.567 0.459 8.803 1.596
X-Restormer [5] 16.568 0.585 0.470 5.989 27.007 18.194 0.607 0.475 7.632 5.910
SFHformer [15] 16.061 0.573 0.471 5.933 23.737 17.995 0.605 0.476 7.010 5.438
MambaIRv2 [13] 17.036 0.602 0.463 6.400 34.987 18.017 0.638 0.465 7.292 54.093
Appendix G Limitations
Although our SurgClean dataset systematically collects real-world degraded surgical images and performs detailed data screening and classification, it still presents two potential limitations. First, First, the scale of our dataset has further space for expansion. SurgClean consists of 134 real surgical cases from different laparoscopic devices in our partner hospital, but the overall size is still relatively small and may not adequately cover all the degradation patterns, anatomical representations, and variations in the intraoperative environment. We will collect more types of surgical cases in future work to advance the development of surgical image degradation techniques. Furthermore, we will enrich the diversity of surgeries, e.g., thoracoscopic and ophthalmic surgeries, to provide different visual feature distributions and degradation patterns. Second, the current surgical samples are collected from a single clinical site, which may cause bias related to specific surgical modalities, instruments, and imaging systems. We will strive to construct multi-center data to promote the generalizability of degradation models for application to a wider range of clinical settings.